OpenStudy (anonymous):

Two rectangles are similar. One has a length of 14 cm and a width of 9 cm, and the other has a width of 3 cm. Find the length of the second rectangle. Round to the nearest tenth if necessary.

OpenStudy (anonymous):

Because they are similar you may infer that their corresponding sides are proportional: \[3/9 = x/14. \]

OpenStudy (anonymous):

They are porportional. 3/9=x/14. Do cross multiplication and the answer would be x=4.67

OpenStudy (anonymous):

A suggestion, offer hints then, let the inquirer solve it. They'll learn more...:)

OpenStudy (anonymous):

Ok...Sorry

OpenStudy (anonymous):

Apology unnecessary...:)